For the differential equation y" + 4y' + 13y = 0, a general solution is of the form y = e-2x(C1sin 3x + C2cos 3x), where C1 and C2 are arbitrary constants. 


Applying the initial conditions y(0) = 4 and y'(0) = 2, find the specific solution. 


y = _______ 
Answer #1

y''+4y'+13y=0

find roots first

x^2+4x+13=0

x_{1,\:2}=\frac{-b\pm \sqrt{b^2-4ac}}{2a}

a=1,\:b=4,\:c=13

x_{1,\:2}=\frac{-4\pm \sqrt{4^2-4\cdot \:1\cdot \:13}}{2\cdot \:1}

\:x_{1,\:2}=\frac{-4\pm \:\sqrt{16-52}}{2}

\:x_{1,\:2}=\frac{-4\pm \:\sqrt{-36}}{2}

\:x_{1,\:2}=\frac{-4\pm \:6i}{2}

\:x_{1,\:2}=-2\pm \:3i

for complex roots general solution is

\:y=e^{at}\left(c_1\cos \left(bt\right)+c_2\sin \left(bt\right)\right)

y=e^{-2t}\left(c_1\cos \left(3t\right)+c_2\sin \left(3t\right)\right)....................(1)

.

here y(0)=4

4=e^{-2\cdot \:0}\left(c_1\cos \left(3\cdot \:0\right)+c_2\sin \left(3\cdot \:0\right)\right)

e^{-2\cdot \:0}\left(c_1\cos \left(3\cdot \:0\right)+c_2\sin \left(3\cdot \:0\right)\right)=4

1\cdot \left(\cos \left(3\cdot \:0\right)c_1+\sin \left(3\cdot \:0\right)c_2\right)=4

1\cdot \left(\cos \left(0\right)c_1+\sin \left(0\right)c_2\right)=4

\left(1c_1+0c_2\right)=4

\left(c_1+0\right)=4

{\color{Red} c_1=4}

put it back in equation 1

y=e^{-2t}\left(4\cos \left(3t\right)+c_2\sin \left(3t\right)\right).................(2)

take derivative

y'= \frac{d}{dt}\left(e^{-2t}\left(4\cos \left(3t\right)+c_2\sin \left(3t\right)\right)\right)

y'=\frac{d}{dt}\left(e^{-2t}\right)\left(4\cos \left(3t\right)+c_2\sin \left(3t\right)\right)+\frac{d}{dt}\left(4\cos \left(3t\right)+c_2\sin \left(3t\right)\right)e^{-2t}

y'=\left(-2e^{-2t}\right)\left(4\cos \left(3t\right)+c_2\sin \left(3t\right)\right)+\left(-12\sin \left(3t\right)+c_2\cos \left(3t\right)\cdot \:3\right)e^{-2t}

y'=-2e^{-2t}\left(4\cos \left(3t\right)+c_2\sin \left(3t\right)\right)+e^{-2t}\left(-12\sin \left(3t\right)+3c_2\cos \left(3t\right)\right)

here y'(0)=-2

-2= \:-2e^{-2\cdot 0}\left(4\cos \left(3\cdot 0\right)+c_2\sin \left(3\cdot 0\right)\right)+e^{-2\cdot 0}\left(-12\sin \left(3\cdot 0\right)+3c_2\cos \left(3\cdot 0\right)\right)

-2= \:-2e^0\left(4\cos \left(0\right)+c_2\sin \left(0\right)\right)+e^0\left(-12\sin \left(0\right)+3c_2\cos \left(0\right)\right)

-2= \:-2\cdot 1\left(4\cdot 1+c_2\cdot 0\right)+1\left(-12\cdot 0+3c_2\cdot 1\right)

-2= \:-2\left(4+0\right)+0+3c_2

-2=\:-2\left(4\right)+3c_2

-2=-8+3c_2

-2+8=3c_2

6=3c_2

{\color{Red} c_2=2}

.

.

put this value in equation 2

y=e^{-2t}\left(4\cos \left(3t\right)+c_2\sin \left(3t\right)\right)

{\color{Red} y=e^{-2t}\left(4\cos \left(3t\right)+2\sin \left(3t\right)\right)}..............answer